# Matrix Profile for Time-Series Anomaly Detection: A Reproducible Open-Source Benchmark on TSB-AD

### 07 Apr 2026 v2

@@ 1 Introduction
As a result, dimensional aggregation is not a minor implementation detail, but one of the central algorithmic decisions in multidimensional MP [8, 7]. This report follows the pre-sorting design, which sorts the dimension-wise comparison scores for each subsequence pair before nearest-neighbor selection [8, 7]. For contrast, the alternative post-sorting design applies dimensional ordering only after dimension-wise nearest neighbors have already been fixed. This matters because pre-sorting allows dimensional ordering to influence neighbor selection itself, rather than only reordering evidence after the neighbors have already been chosen. The rationale for adopting pre-sorting, together with a brief contrast against post-sorting, is discussed in Section˜$2.$ The source paper [7] recommends pre-sorting and reports stronger average benchmark performance for the pre-sorting variants it evaluates.
@@ 2.1 Pipeline Overview
The method can be summarized as a four-stage conceptual pipeline. These stages summarize how the input time series is transformed into the final anomaly score vector. Stage 1 computes the dimension-wise pairwise z-normalized Euclidean distances between subsequence pairs. The resulting pairwise distance tensor has shape $n_{\mathrm{sub}}\times n_{\mathrm{sub}}\times d$, and we denote it by $\mathbf{D}$. This tensor is a conceptual object used to describe the computation, not a tensor that is explicitly materialized in memory by the implementation. Stage 2 applies pre-sorting to $\mathbf{D}$. The purpose and effect of this step are explained in Section˜$2.$ Stage 3 performs efficient exclusion-zone-aware kkNN selection on the pre-sorted distances to produce the multidimensional profile tensor $\mathbf{P}\in\mathbb{R}^{n_{\mathrm{sub}}\times d\times k}$, whose axes correspond to subsequence, ordered profile dimension, and neighbor order. The retrieval step is explained in Section˜2.3. Stage 4 reduces $\mathbf{P}$ to a one-dimensional subsequence score vector $s$ and then applies moving-average smoothing to obtain the final anomaly score vector $y$. The reduction and smoothing step is explained in Section˜2.4. The overall data flow can be summarized compactly as
@@ 2.3 Efficient kkNN Retrieval
The algorithm used here addresses that issue by first applying linear-time selection to keep only a reduced candidate set before sorting [1, 4]. Following the presentation in [7], this reduced set has size kmkm. In the implementation, the same idea is parameterized by the explicit exclusion-zone length $\ell_{\mathrm{ex}}=\lfloor 0.5m\rfloor$, so the reduced candidate count remains proportional to kmkm under the default self-join setting. It then sorts only that reduced candidate set rather than the full set of $n_{\mathrm{ref}}$ candidates [7]. When $n_{\mathrm{ref}}$ is much larger than $km\log(km)$, this yields the same overall scaling reported in [7]: $O(n_{\mathrm{query}}n_{\mathrm{ref}}d)$ instead of $O(kn_{\mathrm{query}}n_{\mathrm{ref}}d)$ for brute force or $O(n_{\mathrm{query}}n_{\mathrm{ref}}\log n_{\mathrm{ref}}d)$ for naive sorting. This algorithmic improvement is a key reason the kkNN extension in [7] reduces the asymptotic cost of multidimensional matrix-profile scoring. Algorithm˜1 summarizes the inner $k$-th-neighbor selection kernel used in the submission.
@@ 3.3 Main Benchmark Results
Table˜2 focuses on the submission configuration selected on the tuning split by average VUS-PR. For the benchmark baselines, both the average VUS-PR values and the rank column are recomputed from the released per-dataset VUS-PR tables [5] after adding the MMPAD setting studied here. In the univariate track, MMPAD reaches an average VUS-PR of 0.4100, second only to Sub-PCA at 0.4234. Its average per-dataset rank is 13.9829, which is also the second best among all compared methods. This places MMPAD among the strongest univariate methods in the benchmark comparison even though it does not take the top mean VUS-PR.
In the multivariate track, MMPAD achieves the highest average VUS-PR at 0.3548, ahead of CNN at 0.3130. However, its average per-dataset rank is 11.2833, which is much weaker than its headline score. This contrast between average score and average rank motivates the dataset-level comparison below.
@@
Univariate Sub-PCA 0.4234 14.2229
MMPAD 0.4100 13.9829
KShapeAD 0.4008 15.4871
POLY 0.3898 13.0286
Series2Graph 0.3881 14.0314
MOMENT (FT) 0.3857 14.3357
MOMENT (ZS) 0.3790 15.0314
KMeansAD 0.3664 15.4143
USAD 0.3612 16.7671
Sub-KNN 0.3501 15.3857
Multivariate MMPAD 0.3548 11.2833
CNN 0.3130 8.1667
OmniAnomaly 0.3122 10.9306
PCA 0.3096 10.2417
LSTMAD 0.3066 9.1333
USAD 0.3041 11.1861
AutoEncoder 0.2950 11.4583
KMeansAD 0.2949 9.2278
CBLOF 0.2731 11.9667
MCD 0.2711 12.3611
@@ 3.4 Performance by Dataset Characteristics
Single anomaly +0.1282 0.6335 161 -0.0904 0.3617 47
Multiple anomalies -0.1340 0.4180 189 +0.0885 0.5113 133
Point anomaly +0.2975 0.7143 49 -0.1368 0.0000 13
Sequence anomaly -0.0636 0.4814 322 +0.0418 0.4722 180
Short series +0.0891 0.5943 175 -0.1274 0.2527 91
Long series -0.1160 0.4400 175 +0.2149 0.6966 89
Low anomaly ratio +0.1975 0.7371 175 +0.0575 0.5000 90
High anomaly ratio -0.2243 0.2971 175 +0.0261 0.4444 90
Short anomalies +0.1621 0.6743 175 -0.0394 0.2889 90
Long anomalies -0.1889 0.3600 175 +0.1230 0.6556 90
@@
  • 2–3 dimensions: +0.2595+0.2595.
  • 4–19 dimensions: +0.0750+0.0750.
  • 20–31 dimensions: 0.0806-0.0806.
  • 32–248 dimensions: 0.1725-0.1725.
@@ 3.5 CPU–GPU Comparison
At the aggregate level, the open-source CPU and GPU implementations produce very similar benchmark results. In the univariate track, the average VUS-PR difference is below 0.0001, and the largest absolute difference on any evaluation dataset is 0.0055. The same conclusion holds in the multivariate track. The average VUS-PR is 0.3548 on GPU and 0.3540 on CPU, with a mean absolute per-dataset difference of 0.0024 and a maximum absolute difference of 0.0692. Taken together, these results provide a useful implementation check: the CPU and GPU versions agree closely in aggregate, even though some individual datasets differ more noticeably.
@@ 3.6 Key Takeaways
  • Matrix-profile-based anomaly detection remains highly competitive on TSB-AD. MMPAD is second on the univariate leaderboard by average VUS-PR and first on the multivariate leaderboard by the same metric.
@@ 4 Conclusion
This report documented the MMPAD submission to TSB-AD from method design through benchmark evaluation. MMPAD is second on the univariate leaderboard by average VUS-PR and first on the multivariate leaderboard by the same metric. The dataset-characteristic analysis also highlights two open challenges for future MP-based anomaly detection. In the univariate track, one recurring gap relative to a strong subspace baseline is broader or repeated anomalous behavior, which suggests testing whether compact dictionary representations of time series [9] can augment the current discord pipeline. In the multivariate track, weaker results on point-anomaly, short-anomaly, and higher-dimensional subsets suggest testing preprocessing that makes normal cross-channel structure more explicit, such as PCA whitening before running MMPAD [6].

### 24 Apr 2026 v3

@@ 1 Introduction
As a result, dimensional aggregation is not a minor implementation detail, but one of the central algorithmic decisions in multidimensional MP [8, 7]. This report follows the pre-sorting design, which sorts the dimension-wise comparison scores for each subsequence pair before nearest-neighbor selection [8, 7]. For contrast, the alternative post-sorting design applies dimensional ordering only after dimension-wise nearest neighbors have already been fixed. This matters because pre-sorting allows dimensional ordering to influence neighbor selection itself, rather than only reordering evidence after the neighbors have already been chosen. The rationale for adopting pre-sorting, together with a brief contrast against post-sorting, is discussed in Section˜2.2. The source paper [7] recommends pre-sorting and reports stronger average benchmark performance for the pre-sorting variants it evaluates.
@@ 2.1 Pipeline Overview
The method can be summarized as a four-stage conceptual pipeline. These stages summarize how the input time series is transformed into the final anomaly score vector. Stage 1 computes the dimension-wise pairwise z-normalized Euclidean distances between subsequence pairs. The resulting pairwise distance tensor has shape $n_{\mathrm{sub}}\times n_{\mathrm{sub}}\times d$, and we denote it by $\mathbf{D}$. This tensor is a conceptual object used to describe the computation, not a tensor that is explicitly materialized in memory by the implementation. Stage 2 applies pre-sorting to $\mathbf{D}$. The purpose and effect of this step are explained in Section˜2.2. Stage 3 performs efficient exclusion-zone-aware kkNN selection on the pre-sorted distances to produce the multidimensional profile tensor $\mathbf{P}\in\mathbb{R}^{n_{\mathrm{sub}}\times d\times k}$, whose axes correspond to subsequence, ordered profile dimension, and neighbor order. The retrieval step is explained in Section˜2.3. Stage 4 reduces $\mathbf{P}$ to a one-dimensional subsequence score vector $s$ and then applies moving-average smoothing to obtain the final anomaly score vector $y$. The reduction and smoothing step is explained in Section˜2.4. The overall data flow can be summarized compactly as
@@ 2.3 Efficient kkNN Retrieval
The algorithm used here addresses that issue by first applying linear-time selection to keep only a reduced candidate set before sorting [1, 4]. Following the presentation in [7], this reduced set has size $km$. In the implementation, the same idea is parameterized by the explicit exclusion-zone length $\ell_{\mathrm{ex}}=\lfloor 0.5m\rfloor$, so the reduced candidate count remains proportional to $km$ under the default self-join setting. It then sorts only that reduced candidate set rather than the full set of $n_{\mathrm{ref}}$ candidates [7]. When $n_{\mathrm{ref}}$ is much larger than $km\log(km)$, this yields the same overall scaling reported in [7]: $O(n_{\mathrm{query}}n_{\mathrm{ref}}d)$ instead of $O(kn_{\mathrm{query}}n_{\mathrm{ref}}d)$ for brute force or $O(n_{\mathrm{query}}n_{\mathrm{ref}}\log n_{\mathrm{ref}}d)$ for naive sorting. This algorithmic improvement is a key reason the kkNN extension in [7] reduces the asymptotic cost of multidimensional matrix-profile scoring. Algorithm˜1 summarizes the inner $k$-th-neighbor selection kernel used in the submission.
@@ 3.3 Main Benchmark Results
Table˜2 focuses on the submission configuration selected on the tuning split by average VUS-PR. For the benchmark baselines, both the average VUS-PR values and the rank column are recomputed from the released per-dataset VUS-PR tables [5] after adding the MMPAD setting studied here. In the univariate track, MMPAD reaches an average VUS-PR of 0.4399, ahead of Sub-PCA at 0.4234. Its average per-dataset rank is 12.5943, which is second only to POLY at 12.4971. This places MMPAD at the top of the univariate benchmark comparison by mean VUS-PR, even though POLY remains slightly stronger by average per-dataset rank.
In the multivariate track, MMPAD achieves the highest average VUS-PR at 0.3539, ahead of CNN at 0.3130. Its average per-dataset rank is 10.7889, indicating that these gains are concentrated on part of the benchmark rather than being uniform across all datasets. This contrast between average score and average rank motivates the dataset-level comparison below.
@@
Univariate MMPAD 0.4399 12.5943
Sub-PCA 0.4234 13.7343
KShapeAD 0.4008 14.9029
POLY 0.3898 12.4971
Series2Graph 0.3881 13.5429
MOMENT (FT) 0.3857 13.8300
MOMENT (ZS) 0.3790 14.5229
KMeansAD 0.3664 14.9614
USAD 0.3612 16.2014
Sub-KNN 0.3501 14.9186
Multivariate MMPAD 0.3539 10.7889
CNN 0.3130 7.6778
OmniAnomaly 0.3122 10.3917
PCA 0.3096 9.7972
LSTMAD 0.3066 8.7000
USAD 0.3041 10.6583
AutoEncoder 0.2950 10.9472
KMeansAD 0.2949 8.6611
CBLOF 0.2731 11.4611
MCD 0.2711 11.7944
@@ 3.4 Performance by Dataset Characteristics
Single anomaly +0.1603 0.6398 161 -0.0809 0.3617 47
Multiple anomalies -0.1060 0.4233 189 +0.0840 0.4962 133
Point anomaly +0.2382 0.7143 49 -0.2006 0.0000 13
Sequence anomaly -0.0231 0.4876 322 +0.0409 0.4611 180
Short series +0.1112 0.5943 175 -0.1331 0.2308 91
Long series -0.0782 0.4514 175 +0.2189 0.6966 89
Low anomaly ratio +0.2225 0.7314 175 +0.0658 0.5111 90
High anomaly ratio -0.1896 0.3143 175 +0.0160 0.4111 90
Short anomalies +0.1857 0.6629 175 -0.0438 0.2778 90
Long anomalies -0.1528 0.3829 175 +0.1256 0.6444 90
@@
  • 2–3 dimensions: +0.2517+0.2517.
  • 4–19 dimensions: +0.0798+0.0798.
  • 20–31 dimensions: 0.0801-0.0801.
  • 32–248 dimensions: 0.1724-0.1724.
@@ 3.5 CPU–GPU Comparison
This subsection is included only as an implementation-consistency check between the open-source CPU and GPU versions. At the aggregate level, the two implementations produce very similar benchmark results. In the univariate track, the average VUS-PR difference is below 0.0001, and the largest absolute difference on any evaluation dataset is 0.0055. The same conclusion holds in the multivariate track. The average GPU-CPU difference is only 0.0008, with a mean absolute per-dataset difference of 0.0024 and a maximum absolute difference of 0.0692. Taken together, these results provide a useful implementation check: the CPU and GPU versions agree closely in aggregate, even though some individual datasets differ more noticeably.
@@ 3.6 Key Takeaways
  • Matrix-profile-based anomaly detection remains highly competitive on TSB-AD. MMPAD is first on both the univariate and multivariate leaderboards by average VUS-PR.
@@ 4 Conclusion
This report documented the MMPAD submission to TSB-AD from method design through benchmark evaluation. MMPAD is first on both the univariate and multivariate leaderboards by average VUS-PR. The dataset-characteristic analysis also highlights two open challenges for future MP-based anomaly detection. In the univariate track, one recurring gap relative to a strong subspace baseline is broader or repeated anomalous behavior, which suggests testing whether compact dictionary representations of time series [9] can augment the current discord pipeline. In the multivariate track, weaker results on point-anomaly, short-anomaly, and higher-dimensional subsets suggest testing preprocessing that makes normal cross-channel structure more explicit, such as PCA whitening before running MMPAD [6].